# Convergence of a Finite Volume Scheme for Compactly Heterogeneous Scalar Conservation Laws

### v2

@@ 3.3.1 Compensated compactness
The compensated compactness method and its applications to systems of conservation laws is for instance reviewed in [10, 29]. Modified for our use, the compensated compactness lemma reads as follows.
@@
Assume that $H$ satisfies (CH)–(CVX), ensuring its genuine nonlinearity in the sense of
$\forall(x,s)\in\mathbb{R}^{2},\quad\text{meas}\biggl{(}\left\{u\in\mathbb{R}\;:\;\partial_{u}H(x,u)=s\right\}\biggr{)}=0.$ (3.13)
@@
$\begin{array}[]{cc}S_{1}(u)\coloneq u-k&Q_{1}(x,u)\coloneq H(x,u)-H(x,k)\\[5.0pt] S_{2}(u)\coloneq H(x,u)-H(x,k)&\displaystyle{Q_{2}(x,u)\coloneq\int_{k}^{u}\partial_{u}H(x,\xi)^{2}\mathinner{\mathrm{d}{\xi}}}.\end{array}$ (3.14)
@@
To prove the $\mathbf{H}_{\mathbf{loc}}^{-1}$ compactness required in Lemma 3.11 we will use the following technical result, see [32, 15, 35].
@@
For the compactness analysis, we take inspiration from [22, 23].
@@
$E_{\Delta}(\phi)\coloneq\int_{0}^{+\infty}\int_{\mathbb{R}}S(u_{\Delta})\partial_{t}\phi+Q(w_{\Delta},u_{\Delta})\partial_{x}\phi\;\mathinner{\mathrm{d}{x}}\mathinner{\mathrm{d}{t}},\quad\phi\in\mathbf{C}_{\mathbf{c}}^{\infty}(\mathbb{R}^{+}\times\mathbb{R},\mathbb{R}).$ (3.15)
@@
$\iint_{P_{j+1/2}^{n}}$ $S(u_{\Delta})\partial_{t}\phi+Q(w_{\Delta},u_{\Delta})\partial_{x}\phi\;\mathinner{\mathrm{d}{x}}\mathinner{\mathrm{d}{t}}$
@@
$+\int_{t^{n}}^{t^{n+1}}[\![\sigma S(u_{\Delta})-Q(w_{\Delta},u_{\Delta})]\!]_{y}\;\phi(t,y(t))\mathinner{\mathrm{d}{t}}$
@@
$[\![\sigma S(u_{\Delta})-Q(w_{\Delta},u_{\Delta})]\!]_{y}\coloneq(\sigma S(u_{\Delta})-Q(w_{\Delta},u_{\Delta}))(t,y(t)+)-(\sigma S(u_{\Delta})-Q(w_{\Delta},u_{\Delta}))(t,y(t)-).$
It is relevant to say that $w_{\Delta}$ is continuous along $\{x=y(t)\}$, equal to either $x_{j}$ or $x_{j+1}$, depending on the sign of $\sigma$.
@@
$\begin{array}[]{clr}E_{\Delta}(\phi)&\displaystyle{=-\int_{\mathbb{R}}S(u_{\Delta}(0,x))\phi(0,x)\mathinner{\mathrm{d}{x}}}&\longleftarrow I_{1}(\phi)\\[5.0pt] &\displaystyle{+\sum_{n=1}^{+\infty}\int_{\mathbb{R}}(S(u_{\Delta}(t^{n}-,x))-S(u_{\Delta}(t^{n},x)))\phi(t^{n},x)\mathinner{\mathrm{d}{x}}}&\longleftarrow I_{2}(\phi)\\[5.0pt] &\displaystyle{+\sum_{n=0}^{+\infty}\sum_{j\in\mathbb{Z}}\sum_{y}\int_{t^{n}}^{t^{n+1}}[\![\sigma S(u_{\Delta})-Q(w_{\Delta},u_{\Delta})]\!]_{y}\;\phi(t,y(t))\mathinner{\mathrm{d}{t}}}&\longleftarrow I_{3}(\phi)\\[5.0pt] &\displaystyle{+\sum_{n=0}^{+\infty}\sum_{j\in\mathbb{Z}}\int_{t^{n}}^{t^{n+1}}\biggl{(}Q(x_{j},u_{\Delta}(t,x_{j+1/2}-))-Q(x_{j+1},u_{\Delta}(t,x_{j+1/2}+))\biggr{)}\phi(t,x_{j+1/2})\;\mathinner{\mathrm{d}{t}}}&\longleftarrow I_{4}(\phi).\end{array}$ (3.16)
@@
Assume that $H$ satisfies (CH)–(CVX). Let $T>0$ and $R>0$. Fix $N,K\in\mathbb{N}^{*}$ such that $T\in[t^{N},t^{N+1}\mathclose{[}$ and $R\in I_{K}$. Then, there exists a constant $c_{1}>0$ depending on $T$, $R$, $u_{o}$, and $H$ such that
$\sum_{n=1}^{N}\sum_{|j|\leq K}\int_{I_{j}}|u_{\Delta}(t^{n},x)-u_{\Delta}(t^{n}-,x)|^{2}\mathinner{\mathrm{d}{x}}\leq c_{1}$ (3.17)
and $\sum_{n=0}^{N}\sum_{|j|\leq K}\sum_{y}\int_{t^{n}}^{t^{n+1}}[\![\sigma S(u_{\Delta})-Q(w_{\Delta},u_{\Delta})]\!]_{y}\;\mathinner{\mathrm{d}{t}}\leq c_{1}.$
Proof.  For the purpose of this proof, let us choose $S(u)=\frac{u^{2}}{2}$ and $Q=Q(x,u)$ its entropy flux. Let $(\phi_{k})_{k}$ be a sequence of nonnegative test functions that converges to $\phi\coloneq\mathbf{1}_{[0,T]\times[-R,R]}$. At the limit $k\to+\infty$ in (3.16), we obtain:
$I_{2}(\phi)+I_{3}(\phi)=\int_{|x|\leq R}|u_{\Delta}(0,x)|^{2}\mathinner{\mathrm{d}{x}}-I_{4}(\phi)\leq 2R\|u_{o}\|_{\mathbf{L}^{\infty}}^{2}-I_{4}(\phi).$
@@
$I_{2}(\phi)$ $=\frac{1}{2}\sum_{n=1}^{N}\sum_{|j|\leq K}\int_{I_{j}}u_{\Delta}(t^{n}-,x)^{2}-u_{\Delta}(t^{n},x)^{2}\;\mathinner{\mathrm{d}{x}}$
$=\frac{1}{2}\sum_{n=1}^{N}\sum_{|j|\leq K}\left\{\int_{I_{j}}(u_{\Delta}(t^{n}-,x)-u_{\Delta}(t^{n},x))^{2}+2u_{\Delta}(t^{n},x)(u_{\Delta}(t^{n}-,x)-u_{\Delta}(t^{n},x))\right\}\mathinner{\mathrm{d}{x}}$
$=\frac{1}{2}\sum_{n=1}^{N}\sum_{|j|\leq K}\int_{I_{j}}(u_{\Delta}(t^{n}-,x)-u_{\Delta}(t^{n},x))^{2}\mathinner{\mathrm{d}{x}}+\sum_{n=1}^{N}\sum_{|j|\leq K}u_{j}\underbrace{\int_{I_{j}}(u_{\Delta}(t^{n}-,x)-u_{j}^{n})\mathinner{\mathrm{d}{x}}}_{=0}.$
@@
$\mathcal{R}_{j+1/2}(k,k)\leq\left(2\|\alpha^{\prime}\|_{\mathbf{L}^{\infty}}\sup_{\underset{p\in[\underline{u},\overline{u}]}{x\in\mathbb{R}}}|\partial_{u}H(x,p)|+\sup_{\underset{p\in[\underline{u},\overline{u}]}{x\in\mathbb{R}}}|\partial_{x}H(x,p)|\right)\Delta x.$
Using an approximation argument to pass from Kruzhkov entropies to any entropy, see [22], we obtain
$I_{4}(\phi)\geq-2TR\left(2\|\alpha^{\prime}\|_{\mathbf{L}^{\infty}}\sup_{\underset{p\in[\underline{u},\overline{u}]}{x\in\mathbb{R}}}|\partial_{u}H(x,p)|+\sup_{\underset{p\in[\underline{u},\overline{u}]}{x\in\mathbb{R}}}|\partial_{x}H(x,p)|\right).$
@@
We can convert (3.17) into an estimate on the spatial variation of the approximate solutions, following the arguments of [16, Page 67]. For the sake of clarity, we set $\rho_{j+1/2}^{n,\pm}\coloneq u_{\Delta}(t^{n},x_{j+1/2}\pm)$.
@@
Assume that $H$ satisfies (CH)–(CVX). Let $T>0$ and $R>0$. Fix $N,K\in\mathbb{N}^{*}$ such that $T\in[t^{N},t^{N+1}\mathclose{[}$ and $R\in I_{K}$. Then, there exists a constant $c_{2}>0$ depending on $T$, $R$, $u_{o}$ and $H$ such that
$\sum_{n=0}^{N}\sum_{|j|\leq K}\left\{(\rho_{j+1/2}^{n,-}-\rho_{j+1/2}^{n,+})^{2}+(u_{j}^{n}-\rho_{j+1/2}^{n,-})^{2}+(u_{j+1}^{n}-\rho_{j+1/2}^{n,+})^{2}\right\}\Delta x\leq c_{2}.$ (3.18)
@@
Assume that $H$ satisfies (CH)–(CVX). Let $(S_{i},Q_{i})_{i\in\{1,2\}}$ be the entropy/entropy flux pairs defined in Lemma 3.11. Then for any $i\in\{1,2\}$, the sequence of distributions $\mu_{i}$ defined by
$\mu_{i}(\phi)\coloneq\int_{0}^{+\infty}\int_{\mathbb{R}}S_{i}(u_{\Delta})\partial_{t}\phi+Q_{i}(x,u_{\Delta})\partial_{x}\phi\;\mathinner{\mathrm{d}{x}}\mathinner{\mathrm{d}{t}},\quad\phi\in\mathbf{C}_{\mathbf{c}}^{\infty}(\mathbb{R}^{+}\times\mathbb{R},\mathbb{R})$
@@
Proof.  We follow the proofs of [22, Lemma 5.5] or [23, Lemma 4.5]. To start, let us fix $(S,Q)$ a smooth entropy/entropy flux pair.
@@
$\mu(\phi)=\int_{0}^{+\infty}\int_{\mathbb{R}}(Q(x,u_{\Delta})-Q(w_{\Delta},u_{\Delta}))\partial_{x}\phi\;\mathinner{\mathrm{d}{x}}\mathinner{\mathrm{d}{t}}+E_{\Delta}(\phi),$
@@
The definition of $w_{\Delta}$ and regularity of $H$ ensures that for any $q_{1}\in\mathopen{]}1,2]$, the first term of the right-hand side is strongly compact in $\mathbf{W}^{-1,q_{1}}(\Omega,\mathbb{R})$. We now estimate $E_{\Delta}(\phi)$.
@@
$|E_{\Delta}(\phi)|\leq\biggl{(}\sup_{\underline{u}\leq p\leq\overline{u}}|S(p)|+\sup_{\underset{\underline{u}\leq p\leq\overline{u}}{x\in\mathbb{R}}}|Q(x,p)|\biggr{)}\cdot\|\phi\|_{\mathbf{W}^{1,\infty}},$
@@
$|I_{1}(\phi)|\leq\|S(u_{o})\|_{\mathbf{L}^{1}([-R,R])}\;\|\phi\|_{\mathbf{L}^{\infty}(\Omega)},\quad I_{3}(\phi)\leq c_{1}\|\phi\|_{\mathbf{L}^{\infty}(\Omega)},$
@@
$I_{2,1}(\phi)$ :-$\coloneq\sum_{n=1}^{N}\sum_{|j|\leq K}\int_{I_{j}}(S(u_{\Delta}(t^{n}-,x))-S(u_{j}^{n}))\phi(t^{n},x_{j})\mathinner{\mathrm{d}{x}}:$
$I_{2,2}(\phi)$ :-$\coloneq\sum_{n=1}^{N}\sum_{|j|\leq K}\int_{I_{j}}(S(u_{\Delta}(t^{n}-,x))-S(u_{j}^{n}))(\phi(t^{n},x)-\phi(t^{n},x_{j}))\mathinner{\mathrm{d}{x}}.:$
We handle $I_{2,1}(\phi)$ by introducing an intermediary point $\omega_{j}^{n}$ such that
$S(u_{\Delta}(t^{n}-,x))-S(u_{j}^{n})=S^{\prime}(u_{j}^{n})(u_{\Delta}(t^{n}-,x)-u_{j}^{n})+\frac{S^{\prime\prime}(\omega_{j}^{n})}{2}(u_{\Delta}(t^{n}-,x)-u_{j}^{n})^{2}.$
@@
$|I_{2,1}(\phi)|\leq\frac{c_{1}}{2}\sup_{\underline{u}\leq p\leq\overline{u}}S^{\prime\prime}(p)\cdot\|\phi\|_{\mathbf{L}^{\infty}(\Omega)},$
@@
$|I_{2,2}(\phi)|$ $\leq\|\phi\|_{\mathbf{C}^{0,\alpha}}\Delta x^{\alpha}\sum_{n=1}^{N}\sum_{|j|\leq K}\int_{I_{j}}|S(u_{\Delta}(t^{n}-,x))-S(u_{j}^{n})|\mathinner{\mathrm{d}{x}}$
$\leq\|\phi\|_{\mathbf{C}^{0,\alpha}}\Delta x^{\alpha-1}\left\{\sum_{n=1}^{N}\sum_{|j|\leq K}\left(\int_{I_{j}}|S(u_{\Delta}(t^{n}-,x))-S(u_{j}^{n})|\mathinner{\mathrm{d}{x}}\right)^{2}\right\}^{1/2}\left\{\sum_{n=1}^{N}\sum_{|j|\leq K}\Delta x^{2}\right\}^{1/2}$
$\leq\sqrt{\frac{2TR}{\lambda}}\|\phi\|_{\mathbf{C}^{0,\alpha}}\Delta x^{\alpha-1/2}\left\{\sum_{n=1}^{N}\sum_{|j|\leq K}\int_{I_{j}}|S(u_{\Delta}(t^{n}-,x))-S(u_{j}^{n})|^{2}\mathinner{\mathrm{d}{x}}\right\}^{1/2}$
@@
Finally, we estimate $I_{4}(\phi)$. Let $k\in[\underline{u},\overline{u}]$ and consider $(S_{i},Q_{i})_{i\in\{1,2\}}$ the entropy/entropy flux pairs defined in Lemma 3.11.
Because of the Rankine-Hugoniot condition,
$\left|[\![Q_{1}(w_{\Delta},u_{\Delta})]\!]_{x_{j+1/2}}^{n}\right|=|H(x_{j},k)-H(x_{j+1},k)|,$
@@
$|I_{4,1}(\phi)|\leq 2TR\;\sup_{x\in\mathbb{R}}|\partial_{x}H(x,k)|\cdot\|\phi\|_{\mathbf{L}^{\infty}(\Omega)},$
@@
$\left|[\![Q_{2}(w_{\Delta},u_{\Delta})]\!]_{x_{j+1/2}}^{n}\right|$ $=\int_{\rho_{j+1/2}^{n,+}}^{\rho_{j+1/2}^{n,-}}\partial_{u}H(x_{j},\xi)^{2}\mathinner{\mathrm{d}{\xi}}+\int_{k}^{\rho_{j+1/2}^{n,+}}\partial_{u}H(x_{j},\xi)^{2}-\partial_{u}H(x_{j+1},\xi)^{2}\mathinner{\mathrm{d}{\xi}}$
$\leq\int_{\rho_{j+1/2}^{n,+}}^{\rho_{j+1/2}^{n,-}}\partial_{u}H(x_{j},\xi)^{2}\mathinner{\mathrm{d}{\xi}}+2L(\overline{u}-\underline{u})\sup_{\underset{\underline{u}\leq p\leq\overline{u}}{x\in\mathbb{R}}}|\partial_{xu}^{2}H(x,p)|\Delta x$
$\leq L\underbrace{\int_{\rho^{-}}^{\rho^{+}}|\partial_{u}H(x_{j},\xi)|\mathinner{\mathrm{d}{\xi}}}_{Q_{2,1}}+2L(\overline{u}-\underline{u})\sup_{\underset{\underline{u}\leq p\leq\overline{u}}{x\in\mathbb{R}}}|\partial_{xu}^{2}H(x,p)|\Delta x,$

### v3

@@ 3.3.1 Compensated compactness
The compensated compactness method and its applications to systems of conservation laws is for instance reviewed in [Chen2000, LuBook]. Modified for our use, the compensated compactness lemma reads as follows. This section culminates with Corollary 3.17, stating the compactness result.
@@
Assume (C3)–(CH)–(CVX) hold, ensuring that $H$ is genuinely nonlinear:
$\forall(x,s)\in\mathbb{R}^{2},\quad\text{meas}\bigl(\left\{p\in\mathbb{R}\;:\;\partial_{u}H(x,p)=s\right\}\bigr)=0.$ (3.14)
@@
$\begin{array}[]{cc}S_{1}(u)\coloneq u-k&Q_{1}(x,u)\coloneq H(x,u)-H(x,k)\\[5.0pt] S_{2}(u)\coloneq H(x,u)-H(x,k)&Q_{2}(x,u)\coloneq\int_{k}^{u}\partial_{u}H(x,\xi)^{2}\mathinner{\mathrm{d}{\xi}}.\end{array}$ (3.15)
@@
The main step of the compensated compactness method is to prove that two entropy productions are compact in $\mathbf{H}_{\mathbf{loc}}^{-1}(\mathopen{]}0,+\infty\mathclose{[}\times\mathbb{R},\mathbb{R})$. It can be obtained using the following technical result, see [Murat1981, DCP1987, SerreBook].
@@
In the remainder of the section, we verify that a generic entropy production of the scheme satisfies the assumptions of Lemma 3.13. Let us stress that this step is nontrivial: it relies on the properties of the solutions to Riemann problems with discontinuous flux. For the analysis, we take inspiration from [KMRTriangular, KMRWellBalanced].
@@
$E_{\Delta}(\phi)\coloneq\int_{0}^{+\infty}\int_{\mathbb{R}}S(u_{\Delta})\partial_{t}\phi+Q(\xi_{\Delta},u_{\Delta})\partial_{x}\phi\;\mathinner{\mathrm{d}{x}}\mathinner{\mathrm{d}{t}},\quad\phi\in\mathbf{C}_{\mathbf{c}}^{\infty}(\mathbb{R}^{+}\times\mathbb{R},\mathbb{R}).$ (3.16)
@@
$\iint_{P_{j+1/2}^{n}}$ $S(u_{\Delta})\partial_{t}\phi+Q(\xi_{\Delta},u_{\Delta})\partial_{x}\phi\;\mathinner{\mathrm{d}{x}}\mathinner{\mathrm{d}{t}}$
@@
$+\int_{t^{n}}^{t^{n+1}}[\![\sigma S(u_{\Delta})-Q(\xi_{\Delta},u_{\Delta})]\!]_{y}\;\phi(t,y(t))\mathinner{\mathrm{d}{t}}$
@@
$\sigma S(u_{\Delta})-Q(\xi_{\Delta},u_{\Delta})]\!]_{y}\coloneq(\sigma S(u_{\Delta})-Q(\xi_{\Delta},u_{\Delta}))(t,y(t)+)-(\sigma S(u_{\Delta})-Q(\xi_{\Delta},u_{\Delta}))(t,y(t)-).$
It is relevant to say that $\xi_{\Delta}$ is continuous along $\{x=y(t)\}$, equal to either $x_{j}$ or $x_{j+1}$, depending on the sign of $\sigma$.
@@
$\begin{array}[]{clr}E_{\Delta}(\phi)&\displaystyle{=-\int_{\mathbb{R}}S(u_{\Delta}(0,x))\phi(0,x)\mathinner{\mathrm{d}{x}}}&\longleftarrow I_{1}(\phi)\\[5.0pt] &\displaystyle{+\sum_{n=1}^{+\infty}\int_{\mathbb{R}}(S(u_{\Delta}(t^{n}-,x))-S(u_{\Delta}(t^{n},x)))\phi(t^{n},x)\mathinner{\mathrm{d}{x}}}&\longleftarrow I_{2}(\phi)\\[5.0pt] &\displaystyle{+\sum_{n=0}^{+\infty}\sum_{j\in\mathbb{Z}}\sum_{y}\int_{t^{n}}^{t^{n+1}}[\![\sigma S(u_{\Delta})-Q(\xi_{\Delta},u_{\Delta})]\!]_{y}\;\phi(t,y(t))\mathinner{\mathrm{d}{t}}}&\longleftarrow I_{3}(\phi)\\[5.0pt] &\displaystyle{+\sum_{n=0}^{+\infty}\sum_{j\in\mathbb{Z}}\int_{t^{n}}^{t^{n+1}}\biggl(Q(x_{j},u_{\Delta}(t,x_{j+1/2}-))-Q(x_{j+1},u_{\Delta}(t,x_{j+1/2}+))\biggr)\phi(t,x_{j+1/2})\;\mathinner{\mathrm{d}{t}}}&\longleftarrow I_{4}(\phi).\end{array}$ (3.17)
@@
Assume (C3)–(CH)–(CVX) hold. Let $T>0$ and $R>0$. Fix $N,K\in\mathbb{N}^{*}$ such that $T\in[t^{N},t^{N+1}\mathclose{[}$ and $R\in I_{K}$. Then, there exists $c_{1}>0$ depending on $T$, $R$, $u_{o}$, and $H$ such that
$\sum_{n=1}^{N}\sum_{{\left|j\right|}\leq K}\int_{I_{j}}{\left|u_{\Delta}(t^{n},x)-u_{\Delta}(t^{n}-,x)\right|}^{2}\mathinner{\mathrm{d}{x}}\leq c_{1}$ (3.18)
and $\sum_{n=0}^{N}\sum_{{\left|j\right|}\leq K}\sum_{y}\int_{t^{n}}^{t^{n+1}}[\![\sigma S(u_{\Delta})-Q(\xi_{\Delta},u_{\Delta})]\!]_{y}\;\mathinner{\mathrm{d}{t}}\leq c_{1}.$
Proof.  For the purpose of this proof, let us choose $S(u)\coloneq\frac{u^{2}}{2}$ and $Q$ its entropy flux. Let $(\phi_{k})_{k}$ be a sequence of nonnegative test functions that converges to $\phi\coloneq\mathbf{1}_{[0,T]\times[-R,R]}$. At the limit $k\to+\infty$ in (3.17), we obtain:
$I_{2}(\phi)+I_{3}(\phi)=\int_{{\left|x\right|}\leq R}{\left|u_{\Delta}(0,x)\right|}^{2}\mathinner{\mathrm{d}{x}}-I_{4}(\phi)\leq 2R{\left\|u_{o}\right\|}_{\mathbf{L}^{\infty}(\mathbb{R})}^{2}-I_{4}(\phi).$
@@
$I_{2}(\phi)$ $=\frac{1}{2}\sum_{n=1}^{N}\sum_{{\left|j\right|}\leq K}\int_{I_{j}}u_{\Delta}(t^{n}-,x)^{2}-u_{\Delta}(t^{n},x)^{2}\;\mathinner{\mathrm{d}{x}}$
$=\frac{1}{2}\sum_{n=1}^{N}\sum_{{\left|j\right|}\leq K}\int_{I_{j}}\biggl\{(u_{\Delta}(t^{n}-,x)-u_{\Delta}(t^{n},x))^{2}+2u_{\Delta}(t^{n},x)(u_{\Delta}(t^{n}-,x)-u_{\Delta}(t^{n},x))\biggr\}\mathinner{\mathrm{d}{x}}$
$=\frac{1}{2}\sum_{n=1}^{N}\sum_{{\left|j\right|}\leq K}\int_{I_{j}}(u_{\Delta}(t^{n}-,x)-u_{\Delta}(t^{n},x))^{2}\mathinner{\mathrm{d}{x}}+\sum_{n=1}^{N}\sum_{{\left|j\right|}\leq K}u_{j}\underbrace{\int_{I_{j}}(u_{\Delta}(t^{n}-,x)-u_{j}^{n})\mathinner{\mathrm{d}{x}}}_{=0}.$
@@
$\mathcal{R}_{j+1/2}(k,k)\leq\bigl(2{\left\|\alpha^{\prime}\right\|}_{\mathbf{L}^{\infty}(\mathbb{R})}\sup_{\underset{\underline{u}\leq p\leq\overline{u}}{x\in\mathbb{R}}}{\left|\partial_{u}H(x,p)\right|}+\sup_{\underset{\underline{u}\leq p\leq\overline{u}}{x\in\mathbb{R}}}{\left|\partial_{x}H(x,p)\right|}\bigr)\Delta x.$
Using an approximation argument to pass from Kruzhkov entropies to any entropy, see [KMRTriangular], we obtain
$I_{4}(\phi)\geq-2TR\bigl(2{\left\|\alpha^{\prime}\right\|}_{\mathbf{L}^{\infty}(\mathbb{R})}\sup_{\underset{\underline{u}\leq p\leq\overline{u}}{x\in\mathbb{R}}}{\left|\partial_{u}H(x,p)\right|}+\sup_{\underset{\underline{u}\leq p\leq\overline{u}}{x\in\mathbb{R}}}{\left|\partial_{x}H(x,p)\right|}\bigr).$
@@
We can convert (3.18) into an estimate on the spatial variation of the approximate solutions, following the arguments of [DiPerna1983, Page 67]. For the sake of clarity, we set $\rho_{j+1/2}^{n,\pm}\coloneq u_{\Delta}(t^{n},x_{j+1/2}\pm)$.
@@
Assume (C3)–(CH)–(CVX) hold. Let $T>0$ and $R>0$. Fix $N,K\in\mathbb{N}^{*}$ such that $T\in[t^{N},t^{N+1}\mathclose{[}$ and $R\in I_{K}$. Then, there exists $c_{2}>0$ depending on $T$, $R$, $u_{o}$ and $H$ such that
$\sum_{n=0}^{N}\sum_{{\left|j\right|}\leq K}\left\{(\rho_{j+1/2}^{n,-}-\rho_{j+1/2}^{n,+})^{2}+(u_{j}^{n}-\rho_{j+1/2}^{n,-})^{2}+(u_{j+1}^{n}-\rho_{j+1/2}^{n,+})^{2}\right\}\Delta x\leq c_{2}.$ (3.19)
@@
Assume that $H$ satisfies (C3)–(CH)–(CVX). Let $(S_{i},Q_{i})_{i\in\{1,2\}}$ be the entropy/entropy flux pairs defined in Lemma 3.12. Then for any $i\in\{1,2\}$, the sequence of distributions $\mu_{i}$ defined by
$\mu_{i}(\phi)\coloneq\int_{0}^{+\infty}\int_{\mathbb{R}}S_{i}(u_{\Delta})\partial_{t}\phi+Q_{i}(x,u_{\Delta})\partial_{x}\phi\;\mathinner{\mathrm{d}{x}}\mathinner{\mathrm{d}{t}},\quad\phi\in\mathbf{C}_{\mathbf{c}}^{\infty}(\mathbb{R}^{+}\times\mathbb{R},\mathbb{R}),$
@@
Proof.  We follow the proofs of [KMRTriangular, Lemma 5.5] or [KMRWellBalanced, Lemma 4.5]. To start, let us fix $(S,Q)$ a smooth entropy/entropy flux pair.
@@
$\mu(\phi)=\int_{0}^{+\infty}\int_{\mathbb{R}}(Q(x,u_{\Delta})-Q(\xi_{\Delta},u_{\Delta}))\partial_{x}\phi\;\mathinner{\mathrm{d}{x}}\mathinner{\mathrm{d}{t}}+E_{\Delta}(\phi),$
@@
The definition of $\xi_{\Delta}$ and regularity of $H$ ensures that for any $q_{1}\in\mathopen{]}1,2]$, the first term of the right-hand side is strongly compact in $\mathbf{W}^{-1,q_{1}}(\Omega,\mathbb{R})$. We now estimate $E_{\Delta}(\phi)$.
@@
${\left|E_{\Delta}(\phi)\right|}\leq\bigl(\sup_{\underline{u}\leq p\leq\overline{u}}{\left|S(p)\right|}+\sup_{\underset{\underline{u}\leq p\leq\overline{u}}{x\in\mathbb{R}}}{\left|Q(x,p)\right|}\bigr)\cdot{\left\|\phi\right\|}_{\mathbf{W}^{1,\infty}(\Omega)},$
@@
${\left|I_{1}(\phi)\right|}\leq{\left\|S(u_{o})\right\|}_{\mathbf{L}^{1}([-R,R])}\;{\left\|\phi\right\|}_{\mathbf{L}^{\infty}(\Omega)},\quad I_{3}(\phi)\leq c_{1}{\left\|\phi\right\|}_{\mathbf{L}^{\infty}(\Omega)},$
@@
$I_{2,1}(\phi)$ $\coloneq\sum_{n=1}^{N}\sum_{{\left|j\right|}\leq K}\int_{I_{j}}(S(u_{\Delta}(t^{n}-,x))-S(u_{j}^{n}))\phi(t^{n},x_{j})\mathinner{\mathrm{d}{x}}$
$I_{2,2}(\phi)$ $\coloneq\sum_{n=1}^{N}\sum_{{\left|j\right|}\leq K}\int_{I_{j}}(S(u_{\Delta}(t^{n}-,x))-S(u_{j}^{n}))(\phi(t^{n},x)-\phi(t^{n},x_{j}))\mathinner{\mathrm{d}{x}}.$
We handle $I_{2,1}(\phi)$ by introducing an intermediary point $\theta_{j}^{n}\in\mathbb{R}$ such that
$S(u_{\Delta}(t^{n}-,x))-S(u_{j}^{n})=S^{\prime}(u_{j}^{n})(u_{\Delta}(t^{n}-,x)-u_{j}^{n})+\frac{S^{\prime\prime}(\theta_{j}^{n})}{2}(u_{\Delta}(t^{n}-,x)-u_{j}^{n})^{2}.$
@@
${\left|I_{2,1}(\phi)\right|}\leq\frac{c_{1}}{2}\sup_{\underline{u}\leq p\leq\overline{u}}S^{\prime\prime}(p)\cdot{\left\|\phi\right\|}_{\mathbf{L}^{\infty}(\Omega)},$
@@
${\left|I_{2,2}(\phi)\right|}$ $\leq{\left\|\phi\right\|}_{\mathbf{C}^{0,\alpha}(\Omega)}\Delta x^{\alpha}\sum_{n=1}^{N}\sum_{{\left|j\right|}\leq K}\int_{I_{j}}{\left|S(u_{\Delta}(t^{n}-,x))-S(u_{j}^{n})\right|}\mathinner{\mathrm{d}{x}}$
$\leq{\left\|\phi\right\|}_{\mathbf{C}^{0,\alpha}(\Omega)}\Delta x^{\alpha-1}\left\{\sum_{n=1}^{N}\sum_{{\left|j\right|}\leq K}\left(\int_{I_{j}}{\left|S(u_{\Delta}(t^{n}-,x))-S(u_{j}^{n})\right|}\mathinner{\mathrm{d}{x}}\right)^{2}\right\}^{1/2}\left\{\sum_{n=1}^{N}\sum_{{\left|j\right|}\leq K}\Delta x^{2}\right\}^{1/2}$
$\leq\sqrt{\frac{2TR}{\lambda}}{\left\|\phi\right\|}_{\mathbf{C}^{0,\alpha}(\Omega)}\Delta x^{\alpha-1/2}\left\{\sum_{n=1}^{N}\sum_{{\left|j\right|}\leq K}\int_{I_{j}}{\left|S(u_{\Delta}(t^{n}-,x))-S(u_{j}^{n})\right|}^{2}\mathinner{\mathrm{d}{x}}\right\}^{1/2}$
@@
Finally, we estimate $I_{4}(\phi)$. Let $k\in[\underline{u},\overline{u}]$ and consider $(S_{i},Q_{i})_{i\in\{1,2\}}$ the entropy/entropy flux pairs defined in Lemma 3.12. Because of the Rankine-Hugoniot condition,
${\left|[\![Q_{1}(\xi_{\Delta},u_{\Delta})]\!]_{x_{j+1/2}}^{n}\right|}={\left|H(x_{j},k)-H(x_{j+1},k)\right|},$
@@
${\left|I_{4,1}(\phi)\right|}\leq 2TR\;\sup_{x\in\mathbb{R}}{\left|\partial_{x}H(x,k)\right|}\cdot{\left\|\phi\right\|}_{\mathbf{L}^{\infty}(\Omega)},$
@@
${\left|[\![Q_{2}(\xi_{\Delta},u_{\Delta})]\!]_{x_{j+1/2}}^{n}\right|}$ $=\int_{\rho_{j+1/2}^{n,+}}^{\rho_{j+1/2}^{n,-}}\partial_{u}H(x_{j},\xi)^{2}\mathinner{\mathrm{d}{\xi}}+\int_{k}^{\rho_{j+1/2}^{n,+}}\partial_{u}H(x_{j},\xi)^{2}-\partial_{u}H(x_{j+1},\xi)^{2}\mathinner{\mathrm{d}{\xi}}$
$\leq\int_{\rho_{j+1/2}^{n,+}}^{\rho_{j+1/2}^{n,-}}\partial_{u}H(x_{j},\xi)^{2}\mathinner{\mathrm{d}{\xi}}+2L(\overline{u}-\underline{u})\sup_{\underset{\underline{u}\leq p\leq\overline{u}}{x\in\mathbb{R}}}{\left|\partial_{xu}^{2}H(x,p)\right|}\Delta x$
$\leq L\underbrace{\int_{\rho^{-}}^{\rho^{+}}{\left|\partial_{u}H(x_{j},\xi)\right|}\mathinner{\mathrm{d}{\xi}}}_{Q_{2,1}}+2L(\overline{u}-\underline{u})\sup_{\underset{\underline{u}\leq p\leq\overline{u}}{x\in\mathbb{R}}}{\left|\partial_{xu}^{2}H(x,p)\right|}\Delta x,$